how would u find the answer of m/12+5/6=5/24
Answer by thechamp1011(19)

TO SOLVE FOR M:
(M/12)+(5/6)=(5/24)
(M/12)=(5/24)-(5/6)
(M/12)=(5-20/24)
(M/12)=(-15/24)
M=(-15/24)*12
M=-15/2=7.5